# Global thermodynamic manifold for conservative control of stochastic systems

### v2

@@ Abstract
Optimal control of stochastic systems plays a central role in nonequilibrium physics, with applications in the study of biological molecular motors and the design of single-molecule experiments. While exact analytic solutions to optimization problems are rare, under slow driving conditions, the problem can be reformulated geometrically solely in terms of equilibrium properties. In this framework, minimum-work protocols are geodesics on a thermodynamic manifold whose metric is a generalized friction tensor. Here, we introduce a new foundation for this friction-tensor formalism for conservatively driven systems. Under complete control of the potential energy, a global thermodynamic manifold (on which points are identified with instantaneous energy landscapes) has as its metric a full-control friction tensor. Arbitrary partial-control friction tensors arise naturally as inherited metrics on submanifolds of this global manifold. Leveraging a simple mathematical relationship between system dynamics and the geometry of the global manifold, we derive new expressions for the friction tensor that offer powerful tools for interpretation and computation of friction tensors and minimum-work protocols. Our results elucidate a connection between relaxation and dissipation in slowly driven systems and suggest optimization heuristics. We demonstrate the utility of these developments in three illustrative examples.
@@ I Introduction
These factors pose challenges to uncovering universal principles governing systems of theoretical and practical interest like biological molecular motors operating in the nonequilibrium environment of living cells—often with remarkable speed and efficiency [3]. Optimal control theory is one framework for navigating the complexity of nonequilibrium statistical physics: given a cost function and boundary conditions, optimization selects a set of privileged paths from the uncountable infinity of thermodynamically distinct paths in control-parameter space [4].
Minimum-work protocols (MWPs) minimize the ensemble average of the excess work, the extra energy required to drive a process in finite time. Although MWPs have been analytically derived for some simple systems [5, 6, 7], the problem is typically intractable. Reformulating the optimal-control problem as a system of Hamiltonian equations allows for the application of numerical methods for ordinary differential equations [6]; however, computation rapidly becomes prohibitively expensive for large systems, high-dimensional control vectors, or systems driven near a critical point. Ongoing efforts to address these challenges include methods employing machine learning [8, 9].
@@
In this paper, we establish a foundational geometric structure for the friction-tensor formalism. We define a non-Riemannian global thermodynamic manifold $\mathcal{M}$ whose metric tensor (24) is the generalized friction tensor obtained by assuming arbitrary control over the microstate energies of a given system. We demonstrate how, in the case of conservative control, arbitrary partial-control friction tensors arise naturally as inherited metrics (30) on submanifolds of $\mathcal{M}$. We then derive two decompositions (42, 54) of the full-control friction tensor that facilitate numerical and analytic calculations, provide insight into the dynamical origins of dissipation for slowly driven systems, and suggest optimization heuristics. We provide concrete demonstrations of the framework in three illustrative examples.
@@ II Conventions and assumptions
That is, $\mu$ is the counting measure for discrete $\Omega$ and the Lebesgue measure for continuous $\Omega$. Then for any state functions $g$ and $h$,
$\left(g,h\right)\equiv\int_{\Omega}\mathrm{d}\mu(x)\,g(x)h(x)$ (2)
is the $\ell^{2}$ inner product (dot product) for discrete $\Omega$ or the $L^{2}$ inner product for continuous $\Omega$. In particular,
@@
Throughout, we extend in the straightforward way the definition of the inner product (2) to arguments other than state functions (e.g., $\left(\bm{u},\bm{v}\right)=\bm{u}^{\mathsf{T}}\bm{v}$ for vectors $\bm{u},\bm{v}\in\mathbb{R}^{m}$).
@@
where in (4b) we define the integral kernel $\mathcal{G}_{t}(x,y)$ of $\mathcal{G}_{t}$, i.e., the rate of probability flowing to $x$ from $y$. For discrete $\Omega$, $\mathcal{G}_{t}$ is a rate matrix $\mathbb{W}_{t}$, and this is the master equation. For continuous $\Omega$, $\mathcal{G}_{t}$ is the second-order differential Fokker-Planck operator $\mathcal{L}_{t}$. The notation $[\mathcal{G}_{t}p(t)](x)$ here reflects that $\mathcal{G}_{t}$ acts on the entire function $p(t)$ (just as $\mathbb{W}_{t}$ acts on the entire vector $\bm{p}(t)$).
We assume that the generator $\mathcal{G}_{\bm{u}}$ satisfies detailed balance for all $\bm{u}$:
@@
where $P(x,t|y,0)=\left[\text{e}^{t\mathcal{G}}\right]\!(x,y)$ is the probability that the system is in state $x$ at time $t$ given that it was initialized in state $y$ at time 00.
@@ III Thermodynamics, control, and geometry
Given a set of control parameters $\bm{u}$ and some cost functional $\mathcal{C}$, an optimal control protocol is a curve $\bm{u}^{*}:\left[0,\tau\right]\to\mathbb{R}^{m}$ that minimizes the cost functional subject to some set of constraints. The choice of cost function depends on the context: e.g., one may wish to minimize entropy production or the loss of available work, maximize power output or efficiency, or optimize some combination of these [27, 1]. In this paper, we study protocols that minimize the average work.
@@
For a protocol $V(t)$ over a time interval $[0,\tau]$, the average work (taken over many instantiations of the control protocol) is obtained by integrating $\left\langle\mathcal{P}(t)\right\rangle$ over the protocol:
@@
It is convenient to instead minimize the mean excess work (or irreversible work), defined as the mean work minus the quasistatic work $\left\langle\mathcal{W}\right\rangle^{(\text{QS})}=\lim_{\tau\to\infty}\left\langle\mathcal{W}\right\rangle$. Rescaling time in (12) with $s=t/\tau$ shows that all dependence on $\tau$ comes from the distribution $p(t)$, so the limit $\tau\to\infty$ is taken by setting $p(t)=\pi(t)$. The mean excess work is then
@@
$\langle\mathcal{W}_{\text{ex}}\rangle=-\int_{0}^{\tau}\mathrm{d}t\left(\frac{\mathrm{d}\bm{u}}{\mathrm{d}t},\left\langle\delta\bm{f}(t)\right\rangle\right)$ (14)
@@
Figure 1: Convergence of the exact excess work $\beta\langle\mathcal{W}_{\text{ex}}\rangle$ (solid red curve) to the linear-response approximation $\beta\langle\mathcal{W}_{\text{ex}}\rangle^{(\text{LR})}$ (dashed teal line) as the protocol duration $\tau$ becomes large, for the $(h,J_{\text{NN}},J_{\text{NNN}},K,L)$ protocol described in Sec. VI.3. As $\tau\to 0$, the excess work asymptotically approaches the relative entropy $D(\pi_{0}\|\pi_{\tau})$ between the initial and final equilibrium distributions [10] (horizontal dashed black line).
We assume that the controlled system is prepared at time $t=-\infty$ with some initial value of the control parameters $\bm{u}_{0}$ so that $p(0)=\pi_{\bm{u}_{0}}$. If the system is driven such that it remains close to equilibrium throughout the protocol, dynamic linear-response theory [29] approximates the average deviation of the conjugate forces $\bm{f}$ to the parameters $\bm{u}$ as
@@
for equilibrium time-correlation matrix $\Sigma_{\bm{f}}(t)=\left\langle\delta\bm{f}(t)\delta\bm{f}(0)^{\mathsf{T}}\right\rangle_{\text{eq}}$. The derivative $\mathrm{d}\Sigma_{\bm{f}}/\mathrm{d}t$ characterizes the system response at time $t$ to a perturbation of the control parameters at time 00.
@@
The linear-response (LR) contribution to the excess work is then
@@
The component $\tilde{\zeta}_{ij}(\bm{u})$ of the generalized friction tensor approximates how the energy cost of simultaneously changing the control parameters $u^{i}$ and $u^{j}$ scales with the velocity of the control parameters. While the exact excess work (13,14) depends on the solution $p(t)$ of the time-inhomogeneous equation (4), the linear-response approximation (19) only requires information about the equilibrium dynamics for fixed parameters. The excess work $\langle\mathcal{W}_{\text{ex}}\rangle$ approaches the linear-response approximation $\langle\mathcal{W}_{\text{ex}}\rangle^{(\text{LR})}$ as the protocol duration $\tau$ increases (Fig.  1).
@@ IV Hierarchy of Thermodynamic Manifolds
In this section, we apply the mathematics of submanifolds  [^2] While this section is intended to be self-contained, the essential geometric ideas will be unfamiliar to many readers. For a thorough overview of differential geometry including submanifolds, degenerate metrics, and extrinsic geometry, with notation that will be more familiar to readers versed in general relativity, see [66]. For a modern treatment of Riemannian submanifolds, see [67]. to connect the geometries of distinct control-parameter sets. By constructing a global thermodynamic manifold $\mathcal{M}$ spanned by the state energies, we show that the partial-control manifolds for arbitrary sets $\left\{u^{0},u^{1},\dots,u^{m-1}\right\}$ of control parameters arise naturally as submanifolds (Fig. 2). In conjunction with the decompositions we will derive in Sec. V, this geometric picture offers methods for rapidly computing friction tensors.
@@ IV.1 Global manifold: Full control
In the context of conservative driving, ‘full control’ refers to the ability to manipulate the entire potential-energy landscape. For a system with $N$ discrete states, the state energies can be assembled into a vector $\bm{V}=(V^{0},V^{1},\dots,V^{N-1})$ with $V^{\nu}=V(x_{\nu})$. For a continuous system, one may imagine a continuum of control parameters $V(x)$ parameterized by $x$. The forces conjugate to the state energies are indicator functions
@@
Analogous to (18), we define the full-control friction tensor $\zeta(V)$—which depends on the entire potential-energy landscape $V$—by its integral kernel
@@
$\left\langle\mathcal{P}_{\text{ex}}(t)\right\rangle^{(\text{LR})}=\left(\frac{\mathrm{d}V}{\mathrm{d}t},\zeta\frac{\mathrm{d}V}{\mathrm{d}t}\right)\equiv\left|\left|\frac{\mathrm{d}V}{\mathrm{d}t}\right|\right|_{\zeta}^{2}\ ,$ (26)
@@
Figure 2: Partial-control manifolds (~$(\tilde{\mathcal{M}},\tilde{\zeta})$ (blue) coordinatized by a set of control parameters $\bm{u}$ are submanifolds of the global thermodynamic manifold $(\mathcal{M},\zeta)$, with $\tilde{\zeta}$ inherited from $\zeta$ by Eq. (30). The restriction $V=V(\bm{u})$ of conservative control defines the immersion $\mathcal{F}$ (28). The pushforward $\mathcal{F}_{*}$—that maps vectors on $\tilde{\mathcal{M}}$ to vectors on $\mathcal{M}$ (red arrows)—is related to the conjugate forces $\bm{f}$ (29).
@@ IV.2 Submanifolds: Partial control
:~$\begin{split}\mathcal{F}&:\tilde{\mathcal{M}}\to\mathcal{M}\\ &:\bm{u}\mapsto V(\bm{u})\end{split}$ (28)
@@
(Substituting (24) into this form and applying (22) shows it to indeed be equivalent to the metric defined in Eq. (18).) This is constructed to ensure that geometric invariants (e.g., the lengths of tangent vectors) are equal when computed on $\mathcal{M}$ and on $\tilde{\mathcal{M}}$. In particular, since the squared length of a vector on a thermodynamic manifold corresponds to the linear-response excess power, the preservation of geometric invariants under coordinate transformations means that one computes the same excess power from $\tilde{\zeta}$ and $\bm{u}(t)$ as from $\zeta$ and $V(t)$. Figure 2 summarizes this picture.
@@ IV.3 Riemannian full-control submanifolds
For a discrete-state system with $|\Omega|=n$, full control is possible with a set of $n-1$ parameters due to the invariance of the dynamics under a uniform shift of the state energies (captured in the degeneracy of $\zeta$). We define a control-parameter set $\bm{u}\in\mathbb{R}^{n-1}$ as complete if the potential energy is fully determined (up to a global additive constant) by $\bm{u}$.
Define the Jacobian $\mathbb{J}=\mathcal{F}_{*}^{\mathsf{T}}$ of the transformation from the control parameters $\bm{u}$ to the state energies $\bm{V}$, i.e., $\left[\mathbb{J}\right]^{\phantom{i}\nu}_{i}=-f_{i}(x_{\nu})=\partial V^{\nu}/\partial u^{i}$. Then for probability vector $\bm{p}(t)$,
@@
In App. A, we show that the completeness condition is equivalent to the invertibility of (31), i.e., the distribution $\bm{p}(t)$ is completely determined by the average conjugate forces $\left\langle\bm{f}\right\rangle_{\bm{p}(t)}$. The inverse of (31) is
@@
for right inverse $\mathbb{J}^{\mathsf{R}}=\mathbb{J}^{\mathsf{T}}\left(\mathbb{J}\mathbb{J}^{\mathsf{T}}\right)^{-1}$. The time evolution of the average deviation $\left\langle\delta\bm{f}(t)\right\rangle$ of the conjugate force to a complete set of control parameters is thus self-contained in the sense that
@@
Figure 3: Schematic of the geometric condition on the completeness of a set of control parameters $\bm{u}$. A control parameter set $\bm{u}$ is complete, i.e., able to fully determine state energies up to a constant offset when the space $\Delta_{f}$ of average conjugate forces $\left\langle\bm{f}\right\rangle$ forms an $(n-1)$-simplex in $\mathbb{R}^{n-1}$ (red triangle, right), since this allows a bijective transformation between $\Delta_{\bm{f}}$ and the probability simplex $\Delta_{\bm{p}}$ (green triangle, left).
[…]
Concretely, consider an $N$-spin system, for which there are $n=2^{N}$ states. We show in Sec. VI.3 that the set of all $k$-spin interactions for $k=1,\dots,N$ from the cluster expansion is a complete control set that spans a $(2^{N}-1)$-dimensional Riemannian full-control submanifold.

### v3

@@ Abstract
Optimal control of stochastic systems plays a central role in nonequilibrium physics, with applications in the study of biological molecular motors and the design of single-molecule experiments. While exact analytic solutions to optimization problems are rare, under slow driving conditions, the problem can be reformulated geometrically solely in terms of equilibrium properties. In this framework, minimum-work protocols are geodesics on a thermodynamic manifold whose metric is a generalized friction tensor. Here, we introduce a foundation for this friction-tensor formalism for conservatively driven systems on either discrete- or continuous-state spaces. Under complete control of the potential energy, a global thermodynamic manifold (on which points are identified with instantaneous energy landscapes) has as its metric a full-control friction tensor. Arbitrary partial-control friction tensors arise naturally as inherited metrics on submanifolds of this global manifold. We leverage a simple mathematical relationship between system dynamics and the geometry of the global manifold to derive expressions for the generalized friction and linear-response excess work. We show that the friction tensor, usually defined as an infinite-time integral, may be expressed as a product of straightforwardly computed matrices. We furthermore show that the linear-response excess power is decomposable into relaxation-time-scaled projections of the control velocity onto eigenmodes of the rate matrix or Fokker-Planck operator. This intuitive result provides a systematic means of interpreting otherwise-mysterious efficient control strategies, unveiling how multi-parameter optimal control takes advantage of multi-exponential relaxation dynamics to reduce dissipation relative to single-parameter control. We demonstrate the computational and conceptual advances in three illustrative examples.
@@ I Introduction
These factors pose challenges to uncovering universal principles governing systems of theoretical and practical interest like biological molecular motors operating in the nonequilibrium environment of living cells—often with remarkable speed and efficiency [3]. Optimal control theory is one framework for navigating the complexity of nonequilibrium statistical physics: given a cost function and boundary conditions, optimization selects a set of privileged paths from the uncountable infinity of thermodynamically distinct paths in control-parameter space [4, 5].
Minimum-work protocols (MWPs) minimize the ensemble average of the excess work, the extra energy required to drive a process in finite time. Although MWPs have been analytically derived for some simple systems [6, 7, 8, 9, 10, 11], the problem is typically intractable. Reformulating the optimal-control problem as a system of Hamiltonian equations allows for the application of numerical methods for ordinary differential equations [11]; however, computation rapidly becomes prohibitively expensive for large systems, high-dimensional control vectors, or systems driven near a critical point. Ongoing efforts to address these challenges include methods employing machine learning [12, 13].
@@
In this paper, we establish a foundational geometric structure for the friction-tensor formalism. We define a non-Riemannian global thermodynamic manifold $\mathcal{M}$ whose metric tensor (24) is the generalized friction tensor obtained by assuming arbitrary control over the microstate energies of a given system. We demonstrate how, in the case of conservative control, arbitrary partial-control friction tensors arise naturally as inherited metrics (30) on submanifolds of $\mathcal{M}$. We then derive two decompositions (43, 55) of the full-control friction tensor that facilitate numerical and analytic calculations, provide insight into the dynamical origins of dissipation for slowly driven systems, and suggest optimization heuristics. We provide concrete demonstrations of the framework in three illustrative examples.
@@ II Conventions and assumptions
Formally, $\mu$ is the counting measure for discrete $\Omega$ and the Lebesgue measure for continuous $\Omega$ [28], i.e., the notation above indicates either a sum or an integral depending on whether $\Omega$ is discrete or continuous. Then for any state functions $g$ and $h$, we denote by
$\left(g,h\right)\equiv\int_{\Omega}\mathrm{d}\mu(x)\,g(x)h(x)\$(2)
the Euclidean $\ell^{2}$ inner product (dot product) for discrete $\Omega$ or the Euclidean $L^{2}$ inner product for continuous $\Omega$. In particular,
@@
Throughout, we extend in the straightforward way the definition of the inner product (2) to arguments other than state functions [e.g., $\left(\bm{u},\bm{v}\right)=\bm{u}^{\mathsf{T}}\bm{v}$ for vectors $\bm{u},\bm{v}\in\mathbb{R}^{m}$].
@@
where in (4b) we define the integral kernel $\mathcal{G}_{t}(x,y)$ of $\mathcal{G}_{t}$, i.e., the rate of probability flowing to $x$ from $y$. For discrete $\Omega$, $\mathcal{G}_{t}$ is a rate matrix $\mathbb{W}_{t}$, and this is the master equation. For continuous $\Omega$, $\mathcal{G}_{t}$ is the second-order differential Fokker-Planck operator $\mathcal{L}_{t}$. The notation $\mathcal{G}_{t}p(t)](x)$ here reflects that $\mathcal{G}_{t}$ acts on the entire function $p(t)$ [just as $\mathbb{W}_{t}$ acts on the entire vector $\bm{p}(t)$].
We further assume that the fixed-parameter dynamics are overdamped and reversible, i.e., the generator $\mathcal{G}_{\bm{u}}$ satisfies detailed balance for all $\bm{u}$:
@@
where $P(x,t|y,0)=\left[\text{e}^{t\mathcal{G}}\right]\!(x,y)$ is the probability that the system is in state $x$ at time $t$ given that it was initialized in state $y$ at time 0.
@@ III Thermodynamics, control, and geometry
Given a set of control parameters $\bm{u}$ and some cost functional $\mathcal{C}$, an optimal control protocol is a curve $\bm{u}^{*}:\left[0,\tau\right]\to\mathbb{R}^{m}$ that minimizes the cost functional subject to some set of constraints. The choice of cost function depends on the context: For example, one may wish to minimize entropy production or the loss of available work, maximize power output or efficiency, or optimize some combination of these [32, 1]. In this paper, we study protocols that minimize the average work subject to a set of holonomic constraints $\bm{V}(t)=\bm{V}(\bm{u}(t))$ that define the control parameters.
@@
For a protocol $V(t)$ over a time interval $0,\tau]$, the average work (taken over many instantiations of the control protocol) is obtained by integrating $\left\langle\mathcal{P}(t)\right\rangle$ over the protocol:
@@
It is convenient to instead minimize the mean excess work (or irreversible work), defined as the mean work minus the quasistatic work $\left\langle\mathcal{W}\right\rangle^{(\text{QS})}=\lim_{\tau\to\infty}\left\langle\mathcal{W}\right\rangle$, obtained by taking $p(t)\to\pi(t)$ in (12). One may verify that the quasistatic work so defined is independent of $\tau$ and is equal to the difference in free energy between the initial and final states.
@@
$\langle\mathcal{W}_{\text{ex}}\rangle=-\int_{0}^{\tau}\mathrm{d}t\left(\frac{\mathrm{d}\bm{u}}{\mathrm{d}t},\left\langle\delta\bm{f}(t)\right\rangle\right)\$(14)
@@
Figure 1: Convergence of the exact excess work $\beta\langle\mathcal{W}_{\text{ex}}\rangle$ (solid red curve) to the linear-response approximation $\beta\langle\mathcal{W}_{\text{ex}}\rangle^{(\text{LR})}$ (dashed teal line) as the protocol duration $\tau$ becomes large, for the $(h,J_{\text{NN}},J_{\text{NNN}},K,L)$ protocol described in Sec. VII.3. As $\tau\to 0$, the excess work asymptotically approaches the relative entropy $D(\pi_{0}\|\pi_{\tau})$ between the initial and final equilibrium distributions [14] (horizontal dashed black line).
We assume that the controlled system is prepared at time $t=$-$\infty$ with some initial value of the control parameters $\bm{u}_{0}$ so that $p(0)=\pi_{\bm{u}_{0}}$. If the system is driven such that it remains close to equilibrium throughout the protocol, dynamic linear-response (LR) theory [34] approximates the average deviation of the conjugate forces $\bm{f}$ to the parameters $\bm{u}$ as
@@
for equilibrium time-correlation matrix $\Sigma_{\bm{f}}(t)=\left\langle\delta\bm{f}(t)\delta\bm{f}(0)^{\mathsf{T}}\right\rangle_{\text{eq}}$. The derivative $\mathrm{d}\Sigma_{\bm{f}}/\mathrm{d}t$ characterizes the system response at time $t$ to a perturbation of the control parameters at time 0.
@@
The LR contribution to the excess work is then
@@
The component $\tilde{\zeta}_{ij}(\bm{u})$ of the generalized friction tensor approximates how the energy cost of simultaneously changing the control parameters $u^{i}$ and $u^{j}$ scales with the velocity of the control parameters. While the exact excess work (13,14) depends on the solution $p(t)$ of the time-inhomogeneous equation (4), the linear-response approximation (19) only requires information about the equilibrium dynamics for fixed parameters. The excess work $\langle\mathcal{W}_{\text{ex}}\rangle$ approaches the linear-response approximation $\langle\mathcal{W}_{\text{ex}}\rangle^{(\text{LR})}$ as the protocol duration $\tau$ increases (Fig. 1).
@@ IV Hierarchy of Thermodynamic Manifolds
In this section, we apply the mathematics of submanifolds  [^2] While this section is intended to be self-contained, the essential geometric ideas will be unfamiliar to many readers. For a thorough overview of differential geometry including submanifolds and degenerate metrics with notation that will be familiar to readers versed in general relativity, see [71]. For a more modern treatment of the subject, see [72]. to connect the geometries of distinct control-parameter sets. By constructing a global thermodynamic manifold $\mathcal{M}$ spanned by the state energies, we show that the partial-control manifolds for arbitrary sets $\left\{u^{0},u^{1},\dots,u^{m-1}\right\}$ of control parameters arise naturally as submanifolds (Fig. 2). In conjunction with the decompositions we will derive in Sec. V, this geometric picture offers methods for rapidly computing friction tensors.
@@ IV.1 Global manifold: Full control
In the context of conservative driving, ‘full control’ refers to the ability to arbitrarily reshape the potential-energy landscape $V(x)$ provided that $\text{e}^{-\beta V(x)}$ remains normalizable. Equivalently, this means that the control-parameter set is sufficiently robust to prepare the controlled system in any equilibrium distribution $\pi$.
For a system with $N$ discrete states, the energy landscape is specified by the energy of all $N$ states, which can be assembled into a vector $\bm{V}=(V^{0},V^{1},\dots,V^{N-1})$ with $V^{\nu}=V(x_{\nu})$. For a continuous system, one may imagine a continuum of control parameters $V(x)$ parameterized by $x$. The forces conjugate to the state energies are indicator functions
@@
Analogous to (18), we define the full-control friction tensor $\zeta(V)$—which depends on the entire potential-energy landscape VV—by its integral kernel
@@
[…]
@@
[…]
@@ IV.2 Submanifolds: Partial control
$\begin{split}\mathcal{F}&:\tilde{\mathcal{M}}\to\mathcal{M}\\ &:\bm{u}\mapsto V(\bm{u})\end{split}$ (28)
@@
Substituting (24) into this form and applying (22) shows it to indeed be equivalent to the metric defined in Eq. (18). This is constructed to ensure that geometric invariants (e.g., the lengths of tangent vectors) are equal when computed on $\mathcal{M}$ and on $\tilde{\mathcal{M}}$. In particular, since the squared length of a vector on a thermodynamic manifold corresponds to the linear-response excess power, the preservation of geometric invariants under coordinate transformations means that one computes the same excess power from $\tilde{\zeta}$ and $\bm{u}(t)$ as from $\zeta$ and $V(t)$. Figure 2 summarizes this picture.
@@ IV.3 Riemannian full-control submanifolds
For a discrete-state system with $|\Omega|=n$, full control is possible with a set of n1n-1 parameters due to the invariance of the dynamics under a uniform shift of the state energies (captured in the degeneracy of $\zeta$). We define a control-parameter set $\bm{u}\in\mathbb{R}^{n-1}$ as complete if the potential energy is fully determined (up to a global additive constant) by $\bm{u}$.
This may be formally stated as follows: a control-parameter set $\bm{u}$ is complete if for all $\bm{V}^{\prime}\in\mathbb{R}^{n}$, there exists some $\bm{u}^{\prime}\in\mathbb{R}^{n-1}$ and $c\in\mathbb{R}$ such that
$\bm{V}^{\prime}=\bm{V}(\bm{u}^{\prime})+c\bm{1}\ .$ (31)
This implies a sort of point-wise generalized invertibility between $\bm{u}$ and $\bm{V}$. If, in addition, $\bm{u}^{\prime}$ is unique for each $\bm{V}^{\prime}$, then we may always project the hypersurface defined by $\bm{V}(\bm{u})$ onto a hyperplane, associating to each control vector a unique energy vector in the plane.
Here we consider complete control sets $\bm{u}$ that only affect relative energies and not the overall offset, so $\bm{V}(\bm{u})$ is a hyperplane with constant normal vector $\bm{1}$.
Define the Jacobian $\mathbb{J}=\mathcal{F}_{*}^{\mathsf{T}}$ of the transformation from the control parameters $\bm{u}$ to the state energies $\bm{V}$, i.e., $\left[\mathbb{J}\right]^{\phantom{i}\nu}_{i}=-f_{i}(x_{\nu})=\partial V^{\nu}(\bm{u})/\partial u^{i}$. Then for probability vector $\bm{p}(t)$,
@@
We show in App. A that under the stated conditions, the distribution $\bm{p}(t)$ is completely determined by the average conjugate forces $\left\langle\bm{f}\right\rangle_{\bm{p}(t)}$, with Eq. (32) invertible via
@@
for right inverse $\mathbb{J}^{\mathsf{R}}=\mathbb{J}^{\mathsf{T}}\left(\mathbb{J}\mathbb{J}^{\mathsf{T}}\right)^{-1}$.
The time evolution of the average deviation $\left\langle\delta\bm{f}(t)\right\rangle$ of the conjugate force to a complete set of control parameters is thus self-contained in the sense that
@@
Figure 3: Schematic of the geometric condition on the completeness of a set of control parameters $\bm{u}$. A control parameter set $\bm{u}$ is complete, i.e., able to fully determine state energies up to a constant offset when the space $\Delta_{f}$ of average conjugate forces $\left\langle\bm{f}\right\rangle$ forms an (n1)(n-1)-simplex in $\mathbb{R}^{n-1}$ (red triangle, right), since this allows a bijective transformation between $\Delta_{\bm{f}}$ and the probability simplex $\Delta_{\bm{p}}$ (green triangle, left).
There is an equivalent geometric condition on the completeness of a control-parameter set, illustrated for a three-state system in Fig. 3. Let $\Delta_{\bm{f}}$ be the convex hull of the $n$ points $\left\{\bm{f}(x_{i})\right\}$, i.e., the set of all conjugate-force averages $\left\langle\bm{f}\right\rangle$. For a complete set of control parameters, $\Delta_{f}$ forms an (n1)(n-1)-dimensional simplex in $\mathbb{R}^{n-1}$. Since the space of probability vectors $\bm{p}$ for an $n$-state system defines an (n1)(n-1)-dimensional simplex (the probability simplex $\Delta_{\bm{p}}=\left\{\bm{p}\in\mathbb{R}^{n}\ :\ \bm{1}^{\mathsf{T}}\bm{p}=1\right\}$), there exists a bijective transformation between these spaces. Put simply, the space of conjugate-force averages $\left\langle\bm{f}\right\rangle$ for a complete set of control parameters is the same “size” as the space of probability vectors.
Concretely, consider an $N$-spin system, for which there are $n=2^{N}$ states. We show in Sec. VII.3 that the set of all $k$-spin interactions for $k=1,\dots,N$ from the cluster expansion is a complete control set that spans a $(2^{N}-1)$-dimensional Riemannian full-control submanifold.